OpenStudy (anonymous):

Find the volume of the solid that is obtained when the region under the curve over the interval [5, 24] is revolved about x-axis.

OpenStudy (anonymous):

bakaloria student?

OpenStudy (anonymous):

please see the attached file

OpenStudy (anonymous):

are you doing bakaloria?

OpenStudy (anonymous):

@agrin no, I dont

OpenStudy (anonymous):

find the integral of the function, put x= 5 and put the equation equal 5, and then youll find C

OpenStudy (anonymous):

do you have a sketch of the graph?

OpenStudy (anonymous):

I got this answer 696.39 is it correct answer??

OpenStudy (anonymous):

put the eq equal to 24*

OpenStudy (anonymous):

like "24 = integral of function (put x=5) + C"

OpenStudy (anonymous):

you have to find C

OpenStudy (anonymous):

I have to find the volume

OpenStudy (anonymous):

you can do this by integration integral of pi ((x+3)^)(2./3) between x = 5 and x = 2424

OpenStudy (anonymous):

and I got the answer 696.39 I want to know if this answer is correct or not?

OpenStudy (anonymous):

then B =24 , A=5 and solve the functions, find its integral

OpenStudy (anonymous):

integral of (x+3)^2/3 = { (x + 3)^(5/3)] / 5/3

OpenStudy (anonymous):

= 3/5 (x+3)^(5/3)

OpenStudy (anonymous):

Here is my solution, please see if this is correct or if not then tell me how do I solve this and what formula should we apply in this situation ..

OpenStudy (anonymous):

your integration is incorrect - its what you got * 3/5

OpenStudy (anonymous):